OpenStudy (anonymous):

Using the substitution v=y/x, solve the IVP y*y' = x^3 + (y^2)/x; y(2) = 6

OpenStudy (anonymous):

whats y*

OpenStudy (anonymous):

y multiplied by y prime i.e. y * dy/dx

OpenStudy (anonymous):

y(y' - y/x ) = x^3 y=vx y' = v + xdv/dx so vx(v + xdv/dx -v) = x^3 vx^2dv/dx = x^3 vdv/dx= x vdv=xdx v^2 = x^2 +c

OpenStudy (anonymous):

there you go..